OpenStudy (anonymous):

Suppose 27 is added as a data value to the set of tata: 10,29,15,29,35, and 2. Which statement is true? A.) The Mean Increases by 4. B.) The Mode decreases by 10. C.) The median decreases by 1. D.) Not Here

OpenStudy (anonymous):

Old Mean = (10+29+15+29+35+2)/ New Mean = (10+29+15+29+35+2+27)/7

OpenStudy (anonymous):

old mean = (10+29+15+29+35+2)/6

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

Old median = 29.5 new 27

OpenStudy (anonymous):

Old mode = new mode = 29

OpenStudy (anonymous):

Mode the most frequent value

OpenStudy (anonymous):

1/2 the values under median

OpenStudy (anonymous):

Mean = average

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Mistake! old median is 22 between 15 and 29 New median is 27

OpenStudy (anonymous):

When you have even number of values, the median is between the two values in the middle as you line up from small to big 2,10,15,median,29,29,35 2,10,15,27,29,29,35 here we have odd number of values and 27 is in the middle